Solve algebraically using logarithms:
2^x = 5^(x+1)

Answer by stanbon(48535)

Take the log of both sides to get:
xlog2 = (x+1)log5
(x+1)/x = log2/log5
1+1/x = 0.43...
1/x = -0.569...
x = -1/0.569
x = -1.756...
Cheers,
Stan H.